# Multiscale Causal Learning

### v2

@@ 3 Arguments and Results
Intervening in the causal chain: We now illustrate how it is possible to “intervene” in the causal chain to disconnect high level behaviour from low level. We show this reduces “utility” (def. 2.2.1) for tasks at higher levels, meaning reduced capacity for WPO. We’re interested in understanding how MCAs relate to abstraction layers in a computer. Just as an organ is really just cells, x86 is a classification we ascribe to the product of human behaviour. It is a higher level behaviour of an MCA that includes humans. To understand how we might “intervene” in the causal chain and disconnect higher level behaviours from the system which gave rise to them, we need to abstract away the lower levels. To relate this back to the cell example, we must abstract away individual cells and focus on the consequences of their behaviour. A policy $\pi\in\Pi_{\alpha_{1}}\cap\Pi_{\alpha_{2}}$ implies an abstraction layer in the sense that for every $o\in E_{\pi}$ there exists $f\in P$ s.t. $\bigcap o=f$ (we can substitute $o$ with a declarative program $f$ that has the same extension and thus truth conditions). We use this fact to justify representing policies as abstraction layers, using a function $\mathfrak{f}$ that takes the extension $E_{\pi}$ of a collective’s policy $\pi$, and returns a new vocabulary $\mathfrak{v^{\prime}}=\{f\in P:\exists o\in E_{\pi}(\bigcap o=f)\}$. We can then use a sequence of uninstantiated tasks $\langle\lambda^{0},\lambda^{1}...\lambda^{n}\rangle$ s.t $\lambda^{i+1}\sqsubset\lambda^{i}$ to represent selection pressures that define “correctness” at every level. At each successive level, “correct” gets more specific. Because everything that happens in a higher level is correct according to lower levels, each higher level must further constrain “correct” if it is to exist in any meaningful sense.
@@
Theorem 1.
The greater the utility $\epsilon(\lambda^{i+1}(\mathfrak{v}^{i+1}))$, the weaker the policy $\pi^{i}$ s.t. $\mathfrak{f}(E_{\pi^{i}})=\mathfrak{v}^{i+1}$ must be (intuitively, this just means adaptability at higher levels implies adaptability at lower levels).
Proof of Theorem 1.
@@
$cells$$organs$$human$$abstractions$
@@
When we create an abstraction layer like the ISA of a computer, we create a machine that mimics our more abstract behaviours. However, the ISA is fixed. It does not learn or adapt the way the lower levels of the human MCA does. It has no agency, the way cells do, because it is not adapting in pursuit of a goal. No graceful degradation. Once we have built an ISA, it exists in the environment independently of the human MCA from which it emerged. This interrupts the causal chain.
$human$$ISA$$abstractions$
@@
As a result, there is no adaptive system underpinning the ISA which can adjust policy at lower levels to facilitate goal directed behaviour at higher levels. Intuitively, if an abstraction layer is a window onto the world, then utility measures how “good” that view is for a task. A static abstraction layer is a window that cannot move, while an MCA can change policy at lower levels, to move the window into an ideal location. When we create an ISA we insert a static abstraction layer at an intermediate level, so the vocabulary at the highest levels is limited to only compositions and permutations of those abstractions we have embodied in the ISA. In comparison, an MCA can learn policies at lower levels, to change the vocabulary of higher levels and facilitate multiscale causal learning. We aren’t suggesting lower levels only optimise for utility at higher levels. After all, an increase in utility may increase the computational cost of searching for correct policies. Furthermore, a spatially extended environment constrains the size of vocabularies at all levels [16], because bounded systems can contain only finite information [32]. We mean lower levels learn an inductive bias for higher levels, balancing utility with other factors in service of goals at every scale.
3.1 Precondition for Collective Identity
According to Cao and Yamins [14] tasks are more “difficult” when they have fewer correct policies, and “easier” when they have many correct policies. In our formalism, Cao and Yamins’ position implies $\mathfrak{v}$-task $\alpha$ is “easier” than $\omega$ if $\mid\Pi_{\alpha}\mid>\mid\Pi_{\omega}\mid$. Collectives that exhibit high level goal directed behaviour that is “easier” have more correct high level policies to choose from. If the policy at a lower level $i-1$ changes such that the task at the higher level $i$ becomes “harder”141414i.e. by reducing $\mid\Pi^{i}\mid$ at the higher level $i$., then this could result in their being no correct policy at level $i$. If there are no longer any correct policies at level $i$ after a change in policy at level $i-1$, and if the system is robust and exhibits graceful degradation, then part of a collective may “break off” (loosening the constraint on correct policies, for both the breakaway part and the part left behind). In more formal terms, say we have a collective of cells $\lambda^{i}(\mathfrak{v}^{i})$ that includes a cell $\alpha\sqsubset\lambda^{i}(\mathfrak{v}^{i})$ s.t. $\mathfrak{f}(E_{\pi^{i-1}_{\mathfrak{a}}})=\mathfrak{v}^{i}_{\mathfrak{a}}$. Now assume the policy $\pi^{i-1}_{\mathfrak{a}}$ changes to $\pi^{i-1}_{\mathfrak{b}}$ s.t. $\Pi_{\lambda^{i}(\mathfrak{v}^{i}_{\mathfrak{b}})}=\emptyset$. A subset of the collective may continue functioning as before with a shared policy, but doing so requires discarding a part of the collective which cannot adopt that policy. This is akin to “selective forgetting” [11] to cope with false negatives, but instead of forgetting data inconsistent with an otherwise good policy, we would “discard” a child task $\alpha\sqsubset\lambda^{i}(\mathfrak{v}^{i}_{\mathfrak{b}})$ of $\lambda^{i}(\mathfrak{v}^{i}_{\mathfrak{b}})$ when $\Pi_{\lambda^{i}(\mathfrak{v}^{i}_{\mathfrak{b}})}=\emptyset$, so that we obtain a task $\lambda^{i^{\prime}}(\mathfrak{v}^{i}_{\mathfrak{b}})\sqsubset\lambda^{i}(\mathfrak{v}^{i}_{\mathfrak{b}})$ s.t. $\alpha\not\sqsubset\lambda^{i^{\prime}}(\mathfrak{v}^{i}_{\mathfrak{b}})$ and $\lambda^{i^{\prime}}(\mathfrak{v}^{i}_{\mathfrak{b}})$ has correct policies, meaning $\Pi_{\lambda^{i^{\prime}}(\mathfrak{v}^{i}_{\mathfrak{b}})}\neq\emptyset$. $\alpha$ is now isolated from the informational structure of the collective and so may have an extremely broad set of “correct” policies to choose from.
@@ 3.1 Precondition for Collective Identity
This is analogous to cancer as cells isolated from a collective informational structure, that lose their high level identity, and pursue lower level goals [27, 28]. If $\alpha$ is a cell, it can now revert to primitive transcriptional behaviours and act like cancer.
4 Discussion
We have formalised MCAs and shown the capacity for WPO at high levels implies WPO at low levels. Causal identities are constructed at different scales. Hence we call it multiscale-causal-learning (MCL). This seems to indicate MCA-like designs could result in more efficient and reliable AI in future.
4.1 Dropping Hume’s Guillotine
Given how we have formalised MCAs, the lowest level is all higher levels (or to put it another way, all higher levels are just classifications we ascribe to parts of the behaviour of the whole). Previous work argues learned behaviour must inevitably be motivated by attraction and repulsion from aspects of the environment [11]. By intervening in the causal chain to create a static abstraction layer, we disconnect higher levels from that motivating impetus. We might describe this as “dropping Hume’s Guillotine” to separate is from ought. This relates to the hard problem of consciousness [34]. According to our formalism, all abstractions constructed by living things are causal identities “goal directed” in the sense of being constructed in response to motivating affect. In dropping Hume’s Guilotine, high level human abstractions would become disconnected from the MCA that constructed them. It may be that, to a human or other MCA, all information amounts to statements of “ought” (motivated, value judgements) [15]. Purely representational content (free of any goal directed motive or connotation) may exist only in the static abstraction layers we engineer.
4.2 Future Research
4.2.1 AI-Safety, Distributed and Cyberphysical-Systems
@@
The UK’s Advance Research and Invention Agency recently initiated a program called “Safeguarded AI” [35]; an ambitious research project that aims to erect a cyberphysical system meeting the standards laid out in Guaranteed Safe AI (GS) [36]. This involves constructing advanced general AI of level ASL-4 [37] [^15] Meaning an artificial general intelligence [38] capable of autonomous research and development that can evaluate the outcomes of deploying less advanced AI, and safeguard humanity against “rogue” AI. in a cyberphysical system. That system seems intended to safeguard against threats from both within and without. The idea of a “rogue AI” from within this system is analogues to cancer, adopting an “unsafe” policy outside the scope of what is permitted by collective. This suggests we could formalise such a system as an MCA, in order to prove properties relevant to AI safety. Control of cyberphysical and distributed systems in general might be framed in terms of an MCA, in order to better understand how different levels in a hierarchy or different parts of a horizontal system constrain each-other.
4.2.2 Active Inference
The sample efficiency of organoids [8] reveals what active inference misses; WPO. Future research might integrate WPO with active inference. Surprise minimisation [7] could be considered an extremely weak policy in and of itself. However, what really matters is fitness [15], not surprise. In adopting machine learning as its foundation, active inference may have inadvertently dropped Hume’s Guillotine, emphasising the wrong goal.

### v3

@@ 3.2 The Causal Chain of Abstraction
A policy $\pi\in\Pi_{\alpha_{1}}\cap\Pi_{\alpha_{2}}$ implies an abstraction layer in the sense that for every $o\in E_{\pi}$ there exists $f\in P$ s.t. $\bigcap o=f$ (we can substitute $o$ with a declarative program $f$ that has the same extension and thus truth conditions). We use this fact to justify representing policies as abstraction layers, using a function $\mathfrak{f}$ that takes the extension $E_{\pi}$ of a collective’s policy $\pi$, and returns a new vocabulary $\mathfrak{v^{\prime}}=\{f\in P:\exists o\in E_{\pi}(\bigcap o=f)\}$. We can then use a sequence of uninstantiated tasks $\langle\lambda^{0},\lambda^{1}...\lambda^{n}\rangle$ s.t $\lambda^{i+1}\sqsubset\lambda^{i}$ to represent selection pressures that define “correctness” at every level. At each successive level, “correct” gets more specific. Because everything that happens in a higher level is correct according to lower levels, each higher level must more tightly constrain what is considered to be “correct” behaviour if that higher level is to exist in a meaningful sense.
@@
Theorem 3 shows that the greater the utility $\epsilon(\lambda^{i+1}(\mathfrak{v}^{i+1}))$, the weaker the policy $\pi^{i}$ s.t. $\mathfrak{f}(E_{\pi^{i}})=\mathfrak{v}^{i+1}$ must be. Intuitively, this just means adaptability later in the cause chain implies adaptability earlier in the chain.
@@
$cells$$humans$$abstractions$
@@
When we create an abstraction layer like the ISA of a computer, we create a machine that embodies a narrow subset of our more abstract behaviours. On top of that ISA it can mimic all behaviours at higher levels of abstraction, because it is Turing complete. However, the ISA itself is fixed. It does not learn or adapt the way the lower levels of the human MCA does. It has no agency, the way cells do, because it is not adapting in pursuit of a goal. No graceful degradation. Once we have built an ISA, it exists in the environment independently of the human MCA from which it emerged. This interrupts the causal chain and creates inefficiencies (which is why, for example, there is demand for application specific chipsets and field programmable gate arrays).
$humans$$ISA$$abstractions$
@@
When we say the ISA is static, we mean there is no adaptive system underpinning the ISA which can adjust policy at lower levels to facilitate goal directed behaviour at higher levels. Intuitively, if an abstraction layer is a window onto the world, then utility measures how “good” that view is for a task. A static abstraction layer is a window that cannot move, while an MCA can change policy at lower levels, to move the window into an ideal location. When we create an ISA we insert a static abstraction layer at an intermediate level, so the vocabulary at the highest levels is limited to only compositions and permutations of those abstractions we have embodied in the ISA. In comparison, an MCA can learn policies at lower levels, to change the vocabulary of higher levels and facilitate multilayer causal learning. What proof 3 demonstrates is that adaptability at higher levels of abstraction demands adaptability at lower levels. We aren’t suggesting lower levels only optimise for utility at higher levels. After all, an increase in utility may increase the computational cost of searching for correct policies. Furthermore, a spatially extended environment constrains the size of vocabularies at all levels [31], because bounded systems can contain only finite information [54]. We mean lower levels learn an inductive bias for higher levels, balancing utility with other factors in service of goals at every scale.
4 Drawbacks of Self Organisation
We have proposed and explored some advantages of multilayer-causal-learning. Are there disadvantages? Do such systems also inherit the flaws of biological systems?
@@ 4 Drawbacks of Self Organisation
Yes, they inherit a state analogous to cancer. In a biological system, when cells are isolated from the informational structure of their collective (i.e. an organ), they lose their high level identity. When that happens, they independently pursue their own lower level goals [55, 56]. They revert to primitive transcriptional behaviour, while the collective of which they were formerly part continues on without them. This is what is commonly known as cancer.
In the context of our formalism, a state analogous to cancer occurs when a system is ’over-constrained’ by the abstraction layer in which it exists. To illustrate what we mean; the more difficult the task the more constrained the set of correct policies [42], meaning $\mathfrak{v}$-task $\alpha$ is easier than $\omega$ if $\mid\Pi_{\alpha}\mid>\mid\Pi_{\omega}\mid$. Collectives that exhibit ‘easier’ high level goal directed behaviour in that sense have more correct high level policies to choose from. They are less tightly constrained. If the policy at a lower level $i-1$ changes such that the task at the higher level $i$ becomes “harder” in the sense of there existing fewer correct policies181818i.e. by reducing $\mid\Pi^{i}\mid$ at the higher level $i$., then this could result in their being no correct policy at level $i$. If there are no longer any correct policies at level $i$ after a change in policy at level $i-1$, then the only way the system can continue to function is for part of the collective to “splinter” or break away from the group, because that will loosen the constraint on correct policies for both the breakaway part and the collective it leaves behind. In more formal terms, say we have a collective of cells $\lambda^{i}(\mathfrak{v}^{i})$ that includes a cell $\alpha\sqsubset\lambda^{i}(\mathfrak{v}^{i})$ s.t. $\mathfrak{f}(E_{\pi^{i-1}_{\mathfrak{a}}})=\mathfrak{v}^{i}_{\mathfrak{a}}$. Now assume the policy $\pi^{i-1}_{\mathfrak{a}}$ changes to $\pi^{i-1}_{\mathfrak{b}}$ s.t. $\Pi_{\lambda^{i}(\mathfrak{v}^{i}_{\mathfrak{b}})}=\emptyset$. A subset of the collective may continue functioning as before with a shared policy, but doing so requires discarding a part of the collective which cannot adopt that policy. This is akin to “selective forgetting” [16] of possible false negatives or positives in order to interpolate otherwise noisy data, but instead of forgetting data inconsistent with an otherwise good policy, we would discard a child task $\alpha\sqsubset\lambda^{i}(\mathfrak{v}^{i}_{\mathfrak{b}})$ of $\lambda^{i}(\mathfrak{v}^{i}_{\mathfrak{b}})$ when $\Pi_{\lambda^{i}(\mathfrak{v}^{i}_{\mathfrak{b}})}=\emptyset$, so that we obtain a task $\lambda^{i^{\prime}}(\mathfrak{v}^{i}_{\mathfrak{b}})\sqsubset\lambda^{i}(\mathfrak{v}^{i}_{\mathfrak{b}})$ s.t. $\alpha\not\sqsubset\lambda^{i^{\prime}}(\mathfrak{v}^{i}_{\mathfrak{b}})$ and $\lambda^{i^{\prime}}(\mathfrak{v}^{i}_{\mathfrak{b}})$ has correct policies, meaning $\Pi_{\lambda^{i^{\prime}}(\mathfrak{v}^{i}_{\mathfrak{b}})}\neq\emptyset$. $\alpha$ is now isolated from the informational structure of the collective and so may have an extremely broad set of “correct” policies to choose from.
This is analogous to cancer. $\alpha$ is isolated and can now pursue goals which do not align with those of the collective. This suggests systems should be under-constrained to ensure graceful degradation or “sloppy fitness” [57], so it can maintain core functionality in as wide a variety of unanticipated circumstances as possible. An under-constrained system may be flexible enough to cope with unexpected inputs that more tightly restrict the set of correct policies. For example, if environment conditions are hostile or “difficult”, that would constitute an input that constrains the space of correct policies. The more difficult the environment, the more tightly constrained behaviour becomes until the system becomes over-constrained and splinters. Conversely, if control of a system is distributed among its parts at the lowest levels (i.e. a swarm rather than a central controller), then the destruction of its parts should not inhibit higher level function. There would be little reason for the parts of a system to splinter, as that would not greatly increase the number of available correct policies.
5 Conclusions and Future Research
We have formalised MCAs as a causal chain, and shown that the capacity for WPO later in the chain depends upon WPO earlier in the chain. This seems to indicate MCA-like designs could result in more adaptable and reliable AI in future. Indeed, the layered structure of artificial neural networks resembles exactly the sort of structure we have described, albeit constrained by the static abstraction layers in which an artificial neural network runs and without WPO. This may explain something of their relative success as a machine learning method. Future research might consider how hardware could be designed in a more distributed, multiscale fashion in order to facilitate MCL by fine tuning abstraction layers. In the context of the current hardware, adaptive abstraction layers might be considered, to facilitate greater efficiency.
Beyond hardware, cyberphysical systems in general could be viewed through the lenses of abstraction and distribution we have described here, to ensure they are robust. We might consider how states analogous to cancer could arise in human organisational structures, or how power may be distributed throughout an organisation to ensure MCL. Free-market trade is a great example of how loose constraints have already facilitated adaptability in large collectives, and which is already being explored in the broader context of MCAs [58]. Likewise, the delegation of greater control has been shown to improve the effectiveness of military organisations [59]. Our claim boils down to a principle of design like an epistemological razor [34]; systems should be no more constrained than absolutely necessary.
@@ 5 Conclusions and Future Research
One concrete example of how this might be applied in future is the regulation of AI. The UK’s Advance Research and Invention Agency recently initiated a program called “Safeguarded AI” [60]; an ambitious research project that aims to erect a cyberphysical system meeting the standards laid out in Guaranteed Safe AI (GS) [61]. This involves constructing advanced general AI of level ASL-4 [62] [^19] Meaning an artificial general intelligence [20] capable of autonomous research and development that can evaluate the outcomes of deploying less advanced AI, and safeguard humanity against “rogue” AI. in a cyberphysical system. That system seems intended to safeguard against threats from both within and without. The idea of a “rogue AI” from within this system is analogous to cancer, adopting an “unsafe” policy at odds with that of the collective. However if we view this system as MCA, then tight regulations may over-constrain the system trigger the cancerous state they are meant to prevent.
This also might be applied to problems of a philosophical nature. Given how we have formalised MCAs, the lowest level is all higher levels (or to put it another way, all higher levels are just classifications we ascribe to parts of the behaviour of the whole). Previous work argues learned behaviour must inevitably be motivated by attraction and repulsion from aspects of the environment [16]. By intervening in the causal chain to create a static abstraction layer, we disconnect higher levels from that motivating impetus. We might describe this as “dropping Hume’s Guillotine” to separate is from ought. This relates to the hard problem of consciousness [63]. All abstractions constructed by living things are “goal directed” in the sense that they are consequence of an evolutionary process [30]. In dropping Hume’s Guilotine, high level human abstractions would become disconnected from the MCA that constructed them. It may be that, to a human or other MCA, all information amounts to statements of “ought” (motivated, value judgements). Purely representational content (free of any goal directed motive or connotation) may exist only in the static abstraction layers we engineer.
In summary, we have formally described biological self organisation across distribution of scale and levels of abstraction, and shown adaptability at higher levels of abstraction depends upon adaptability at lower levels. We have described how a failure state analogous to cancer can occur when a system is too tightly constrained by the abstraction layer in which it exists (i.e. the task it faces is impossible). Though our result is of a purely theoretical nature, it illustrates how we may design more robust systems, and lays a foundation for future empirical research.